OpenStudy (anonymous):

Multiply. (5x2 + x – 4)(x + 2)

OpenStudy (anonymous):

(5x^2 + x – 4)(x + 2) = (x + 2) (5x^2 + x – 4) = x (5x^2 + x – 4) + 2 (5x^2 + x – 4) = 5x^3 + x^2 - 4x + 10x^2 + 2x - 8 = 5x^3 + 11x^2 - 2x - 8